Consider the function f(n) = 35n3+ 2n3log(n) − 2n2log(n2) which represents the complexity of some algorithm.

(a) Find a tight big-O bound of the form g(n) = np for the given function f with some natural number p. What are the constants C and k from the big-O definition?

(b) Find a tight big-Ω bound of the form g(n) = np for the given function f with some natural number p. What are the constants C and k from the big- Ω definition?

(c) Can we conclude that f is big−Θ (np) for some natural number p?



1
Expert's answer
2021-10-06T16:32:26-0400

a)

"f(n)=O(g(n))" if

"Cg(n)\ge f(n)"

"C>0,\ n\ge n_0"


So:

for "n_0=1"

"36n^3\ge 35n^3+ 2n^3log(n) − 2n^2log(n^2)"

"C=36,\ p=3,\ k=1"


b)

"f(n)=O(g(n))" if

"Cg(n)\le f(n)"

for "n_0=1"

"35n^3\le 35n^3+ 2n^3log(n) − 2n^2log(n^2)"

"C=35,\ p=3,\ k=1"


c)

"f(n)=\theta(g(n))" if

"C_2g(n)\le f(n)\le C_1g(n)"


for "n_0=1"

"35n^3\le 35n^3+ 2n^3log(n) − 2n^2log(n^2)\le 36n^3"

So, we can conclude that "f(n)=\theta(g(n))" for "p=3"